OpenStudy (cinniomon):

I still need help on this equation 3y-2(y-1)/6=-1 all of the equation is over 6.

OpenStudy (anonymous):

y=-6

OpenStudy (radar):

3y-2(y-1) --------- =-1 6 clear parenthesis 3y-2y+2 --------- =-1 6 combine similar terms y+2 ----- =-1 6 Multiply both sides by 6 y+2=-6 add -2 to both sides. y=-8

OpenStudy (amistre64):

if you still have questions, this is the place you would ask them to get clarification. having this personal interaction with those who are willing to help you out....

OpenStudy (cinniomon):

well, do I multiply the 2 by 6 as well?

OpenStudy (amistre64):

id say no; the multiplying by 6 creates a condition such that 6/6 = 1

OpenStudy (amistre64):

also known as "canceling" out, or clearing the fraction

OpenStudy (cinniomon):

so it's 6y+2=-6?

OpenStudy (amistre64):

\[\frac{y+2}{6}=-1\] \[\frac{6(y+2)}{6}=-1(6)\] \[\frac{6}{6}(y+2)=-6\] \[1(y+2)=-6\] \[y+2=-6\]

OpenStudy (amistre64):

does that make sense?

OpenStudy (cinniomon):

Thanks a lot! Makes much more sense! :D